OpenStudy (anonymous):

Finding angle between vector F and S. F= (4,-4) and S=(4,3). I tried doing Cosθ= f⋅s/ lFl x lSl I ended up getting θ=90? but thats because absolute value of F ended as 0 ( sqrt[(4^2)+(-4^2)]= 0 but i don't think thats right.. lol help please?

jimthompson5910 (jim_thompson5910):

|F| = sqrt( 4^2 + 4^2 ) |F| = sqrt( 16 + 16 ) |F| = sqrt( 2*16 ) |F| = sqrt( 16*2 ) |F| = sqrt(16)*sqrt(2) |F| = 4*sqrt(2) Not sure how you got |F| = 0 which implies that vector F is 0 units long

jimthompson5910 (jim_thompson5910):

|S| = sqrt(4^2 + 3^2) |S| = sqrt(16+9) |S| = sqrt(25) |S| = 5

jimthompson5910 (jim_thompson5910):

F dot S = <4, 4> dot <4,3> F dot S = 4*4 + 4*3 F dot S = 16+12 F dot S = 28

jimthompson5910 (jim_thompson5910):

do you see how to finish up?

OpenStudy (anonymous):

Sorry ! I had a typo, my Fy=-4 not 4 sorry i just corrected it

jimthompson5910 (jim_thompson5910):

Not much will change |F| = sqrt( 4^2 + (-4)^2 ) |F| = sqrt( 16 + 16 ) |F| = sqrt( 2*16 ) |F| = sqrt( 16*2 ) |F| = sqrt(16)*sqrt(2) |F| = 4*sqrt(2)

jimthompson5910 (jim_thompson5910):

F dot S = <4, -4> dot <4,3> F dot S = 4*4 + (-4)*3 F dot S = 16-12 F dot S = 4

OpenStudy (anonymous):

Ah. ok I see what I did, I input my equation into the calculator wrong :( thank you very much!

jimthompson5910 (jim_thompson5910):

np